# Predictive Coding Networks and Inference Learning: Tutorial and Survey

### v1

@@ 2.1.1. Training Procedure
FNNs are defined by a set of activation nodes organized in a hierarchy of $L$ layers: $\bm{a}^{\ell}\in\mathbb{R}^{n_{\ell}+1}$, where $\ell=0,...,L$ is the layer and $n_{\ell}$ the number of nodes in that layer (a constant value of 1 is added to account for the bias). Layers are connected by a set of weight matrices $\bm{w}^{\ell}\in\mathbb{R}^{n_{\ell}\times(n_{\ell+1}+1)}$ in the following way (bishop_pattern_2006, ; goodfellow_deep_2016, ):
@@
where throughout, the biases $\bm{b}^{\ell}$ are absorbed as an additional column in the weight matrix (as is commonly done (goodfellow_deep_2016, )). We will call this equation the activity rule since it states how activation nodes are computed. The function $f:\mathbb{R}^{n_{\ell}}\rightarrow\mathbb{R}^{n_{\ell}}$ here is a nonlinear activation function (e.g. sigmoid, ReLU). Setting the bottom layer $\bm{a}^{0}$ to a datapoint $\bm{x}^{(n)}$, the final layer $\bm{a}^{L}$ is taken as the predicted label $\hat{\bm{y}}=\hat{\bm{y}}(\bm{w},\bm{x}^{(n)})$. We also define the argument of $f$ as the pre-activation $\bm{z}^{\ell}=\bm{w}^{\ell-1}\bm{a}^{\ell-1}$. Then, notice that the neural network is simply a function $\operatorname{NN}:X\rightarrow Y$, such that $\hat{\bm{y}}=\operatorname{NN}(\bm{x}^{(n)})$ and
@@
(3) $\bm{w}^{\ell}=\underset{\bm{w}}{\operatorname{argmin}}\,\mathcal{L}(\bm{w})\implies\Delta\bm{w}^{\ell}=-\alpha\frac{\partial\mathcal{L}}{\partial\bm{w}^{\ell}},$
where the weights are updated iteratively using $\Delta\bm{w}^{\ell}$, where $\alpha$ is the learning rate. For better training efficiency, in practice one uses stochastic gradient descent (SGD), often with optimizers such as Adam (kingma_adam_2017, ), though this will not be important for our purposes.
@@ 2.1.2. Learning Algorithm
(5) $\Delta\bm{w}^{\ell}=\alpha\bm{\delta}^{\ell+1}\odot f^{\prime}(\bm{w}^{\ell}\bm{a}^{\ell})(\bm{a}^{\ell})^{T}$
@@
It can be observed how errors are computed from the final to the first layer, explaining the name backpropagation of errors. Pseudocode is shown in alg. 1.
@@ 2.1.4. Testing Procedure
Finally, after a network has been trained, testing the network is done by using the activity rule (1) to obtain predicted outputs $\hat{\bm{y}}$ from datapoints $\bm{x}^{(n)}$.
@@ 2.2.1. Direction in PCNs
Before explaining PCNs in detail we briefly mention a convention issue. The literature contains conflicting notions of direction in PCNs, which can cause confusion and are worth discussing at the outset. There are two connected issues. The first is mathematical: should local predictions $\bm{\mu}^{\ell}$ be defined as $\bm{\mu}^{\ell}=f(\bm{w}^{\ell-1}\bm{a}^{\ell-1})$ (as in this section) or $\bm{\mu}^{\ell}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$ (implied by fig. 1)? The second issue is terminological: how should the words forward/backward be used?
Typically in the neuroscientific PC literature (rao_predictive_1999, ; friston_hierarchical_2008, ), predictions are $\bm{\mu}^{\ell}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$, and data is clamped to $\ell=0$ (fig. 1). This means that predictions go towards data (and errors go away from data). In this ‘PC convention’, the word forward is typically chosen to mean away from the data, which is the direction of errors.
@@
However, in the supervised learning context, for a PCN to be equivalent to FNNs during testing, it turns out that one needs to have predictions go away from data. With the conventions above, this means one would have to clamp data to the highest layer $\ell=L$, which is confusing from the perspective of standard ML. This also leads to ambiguity as to whether ‘forward’ should mean towards increasing layer number, or away from the data.
@@
The alternative when comparing PCNs to FNNs is to keep data clamped to $\ell=0$ but change the direction of local prediction to $\bm{\mu}^{\ell}=f(\bm{w}^{\ell-1}\bm{a}^{\ell-1})$, as is done by many recent works (millidge_predictive_2022, ; rosenbaum_relationship_2022, ; salvatori_reverse_2022, ; zahid_predictive_2022, ; millidge_backpropagation_2022, ; alonso_theoretical_2022, ; millidge_theoretical_2022, ; innocenti_understanding_2023, ; song_inferring_2024, ; alonso_understanding_2024, ; byiringiro_robust_2022, ; pinchetti_predictive_2022, ). Forward then means away from the data (equivalently, towards higher layers) as in the PC convention above, except that this is now the direction of predictions. In other words, rather than swapping the data clamping from $\ell=0$ to $L$, we instead swap the directions of predictions and errors relative to fig. 1, so that – when comparing PCNs to FNNs specifically – predictions go forwards and errors go backwards.
@@
The disadvantage of this choice is that $\bm{\mu}^{\ell}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$ with data at $\ell=0$ is remains the more sensible convention for PCNs when seen as Bayesian/generative models (section 3) (rao_predictive_1999, ; salvatori_associative_2021, ; ororbia_neural_2022, ; ofner_generalized_2022, ); it is only in the comparison with FNNs where some inversion (either of clamping or of directions) is required. Hence, in this work we use both conventions depending on the use of the network. We define discriminative PCNs as $\bm{\mu}^{\ell}=f(\bm{w}^{\ell-1}\bm{a}^{\ell-1})$ (for supervised learning), and generative PCNs as $\bm{\mu}^{\ell}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$ (for unsupervised learning; this reflects also the use of these terms in (millidge_predictive_2022-1, ; innocenti_understanding_2023, )). As for terminology, we can then unambiguously use ‘forward’ to mean away from the data in all cases, consistent with standard ML. One must simply bear in mind that for discriminative PCNs, ‘forward’ reflects the direction of predictions, while for generative PCNs it refers to the direction of errors.
@@ 2.2.2. Training Procedure
(not to be confused with the ‘global’ prediction $\hat{\bm{y}}$ of the network as a whole). Note that predictions here go towards lower layers, which is opposite to what is suggested by fig. 1. We discuss this convention in detail in section 2.2.1. The weights and biases are defined just like in FNNs. [^2] Writing out bias, we here use $\bm{\mu}^{\ell}=f(\bm{w}^{\ell-1}\bm{a}^{\ell-1}+\bm{b}^{\ell-1})$, which is the most common convention in ML (goodfellow_deep_2016, ; prince_understanding_2023, ). Most works in the PC literature however, define predictions as $\bm{\mu}^{\ell}=\bm{w}^{\ell-1}f(\bm{a}^{\ell-1})+\bm{b}^{\ell-1}$, whereas the main publicly available implementation uses yet another definition: $\bm{\mu}^{\ell}=f(\bm{w}^{\ell-1}\bm{a}^{\ell-1})+\bm{b}^{\ell-1}$ (tschantz_infer-activelypypc_2023, ). This choice of convention often does not make a difference for performance, but it is worth noting that it does change the form of the update rules. Also, the choice for the last layer is of great practical importance, since it determines the domain of possible outputs (e.g. a sigmoid or softmax is appropriate for classification but fatal for regression). The discrepancy between actual and predicted activity is
@@
often called the energy function. [^3] Note that this energy is the so-called complete data log-likelihood for a particular choice of the generative model, when viewing the PCN as a probabilistic graphical model (section 3). Training now happens by setting the bottom layer $\bm{a}^{0}$ to the datapoint $\bm{x}^{(n)}$, like in FNNs, and additionally fixing the final layer $\bm{a}^{L}$ to the correct label $\bm{y}^{(n)}$. The activity rule is not defined as in (1); instead, the updated values of the hidden (unclamped) activity nodes (i.e. $\ell=1,...,L-1$) are the minimum of the energy function:
@@
This is called the inference phase, since we are inferring the appropriate activation values for hidden nodes, given the clamped nodes (see section 3 for a detailed discussion of the probabilistic interpretation). Note that we use a hat for the optimized values where $E$ is at a minimum.
@@
$=-\gamma\left(\bm{\epsilon}^{\ell}-\left(\bm{w}^{\ell}\right)^{T}\left(\bm{\epsilon}^{\ell+1}\odot f^{\prime}\left(\bm{w}^{\ell}\bm{a}^{\ell}\right)\right)\right)$
@@
again using a hat for the values after minimizing, and a step-wise gradient descent updates which can be done using SGD or Adam. The fact that (11) is trained to convergence before the weight update is performed suggests that training a PCN is more computationally costly than FNNs. However, this added cost can be at least partially circumvented in practice, as we discuss in section 4. Additionally, note that since the energy function is a sum over layers, the update rule does not require backpropagating errors. Instead, taking the derivative in (9) simply gives:
(13) $\Delta\bm{w}^{\ell}=\alpha\bm{\epsilon}^{l+1}\odot f^{\prime}\left(\bm{w}^{\ell}\hat{\bm{a}}^{\ell}\right)(\hat{\bm{a}}^{\ell})^{T},$
i.e., the weight updates can be computed using the error and activation nodes in neighboring layers, meaning PCNs are local both in space and time. We discuss the advantages of this property below.
@@ 2.2.3. Learning Algorithm
By virtue of the use of the energy function for both activation updates and weight updates, training PCNs is not done by backprop, but a different learning algorithm. Pseudocode is shown in alg. 2. This is sometimes called inference learning (IL) (salvatori_associative_2021, ; alonso_theoretical_2022, ; alonso_understanding_2023) or, rarely, prospective configuration (song_inferring_2024, ), but is often simply referred to as ‘predictive coding’ (salvatori_learning_2022, ; salvatori_stable_2023, ). Importantly, it is an instance of expectation maximization (EM) (see section 3). We choose the term IL in this work. Note that unlike BP, IL requires initializing hidden nodes before inference. This is often done using a ‘feedforward pass’, setting $\bm{a}^{\ell}(0)=\bm{\mu}^{\ell}$. This causes the error throughout the network to be zero except at the final layer. Other initializations can also be used (frieder_non-convergence_2022, ).
@@
Importantly, recent work introduced incremental PC (salvatori_stable_2023, ) using the insight that EM admits the use of partial steps, meaning that the minimization of $E$ w.r.t. activations does not have to be complete for the algorithm to converge. Thus, a minor change in alg. 2 gives alg. 3, which we to as incremental IL in this work.
@@ 2.2.4. Structure
A PCN is typically visualized as in fig. 5. The grey connections are the predictions defined by (7), together with error computations, (8) (similar to those in FNNs, see fig. 4). The red connections are defined by (11).
@@ 2.2.5. Testing Procedure
A PCN can be tested by clamping the lowest layer $\bm{a}^{0}$ to the data, and computing updates for all the other layers, now including the final layer $\bm{a}^{L}$, using the activity rule (10). Interestingly, computing these updates is much simpler than during training. With the output nodes now unclamped, (10) has an analytical solution. One can show (whittington_approximation_2017, ; song_predictive_2021, ) (see appendix B.2) that with $\bm{a}^{0}=\bm{x}^{(n)}$, the minimum of $E$ is:
@@
In other words, during testing, a prediction of the network can be computed using a single pass through the network, without requiring gradient-based optimization of activation nodes of (11). In other words, we can even say that testing is equivalent in PCNs and FNNs, or that during testing the PCN ‘becomes’ an FNN in the sense of (2). [^4] Note that ‘testing’ is often more generally called ‘inference’, to refer to use of a network after it has been trained. To prevent confusion with the inference phase during learning in PCNs, we refrain from using this term here. As such, important results for neural networks such as universal approximation theorems (prince_understanding_2023, ) also hold for discriminative PCNs. For these networks, the difference w.r.t. FNNs is only in their learning algorithm. We emphasize that this holds only for discriminative PCNs, with the local prediction defined as in (7), and data clamped to $\bm{a}^{0}$. If this is changed, as will be done in later sections, then this equivalence no longer holds.
@@ 2.3.1. Training & Testing Procedures
table 2 summarizes the important steps and computations for FNNs (with MSE loss) and discriminative PCNs, during training and testing. Notice the equality between the MSE loss and PCN’s energy function during testing, and their similarity during training. This is a result of fixing the final layer of the PCN $\bm{a}^{L}$ the label $\bm{y}^{(n)}$, and labelling the prediction of the last layer $\bm{\mu}^{L}=f(\bm{w}^{L-1}\bm{a}^{L-1})$ as $\hat{\bm{y}}$. The term $\sum_{\ell=1}^{L-1}(\bm{\epsilon}^{\ell})^{2}$ is called the internal error (millidge_backpropagation_2022, ) or residual error (millidge_theoretical_2022, ).
Table 2. Training & testing procedures compared: FNNs vs. discriminative PCNs. Key steps and corresponding equations are shown. For clarity, biases are omitted. A key difference between discriminative PCNs and FNNs is that during training, the former additionally clamps the output to the label, $\bm{a}^{L}=\bm{y}^{(n)}$: this is the fixed output. At the same time, as each layer is predicted by another layer, PCNs also have a predicted output, $\bm{\mu}^{L}$. This corresponds to the activity of the last layer in FNNs, $\bm{a}^{L}$.

### v2

@@ 2.1.1. Training Procedure
FNNs are defined by a set of activation nodes organized in a hierarchy of $L$ layers: $\bm{a}^{\ell}\in\mathbb{R}^{n_{\ell}+1}$, where $\ell=0,...,L$ is the layer and $n_{\ell}$ the number of nodes in that layer (a constant value of 1 is added to account for the bias). Layers are connected by a set of weight matrices $\bm{w}^{\ell}\in\mathbb{R}^{n_{\ell}\times(n_{\ell+1}+1)}$ in the following way (Bishop, 2006; Goodfellow et al., 2016):
@@
where throughout, the biases $\bm{b}^{\ell}$ are absorbed as an additional column in the weight matrix (as is commonly done (Goodfellow et al., 2016)). We will call this equation the activity rule since it states how activation nodes are computed. The function $f:\mathbb{R}^{n_{\ell}}\rightarrow\mathbb{R}^{n_{\ell}}$ here is a nonlinear activation function (e.g. sigmoid, ReLU). Setting the bottom layer $\bm{a}^{0}$ to a datapoint $\bm{x}^{(n)}$, the final layer $\bm{a}^{L}$ is taken as the predicted label $\hat{\bm{y}}=\hat{\bm{y}}(\bm{w},\bm{x}^{(n)})$. Then, notice that the neural network is simply a function $\operatorname{NN}:X\rightarrow Y$, such that $\hat{\bm{y}}=\operatorname{NN}(\bm{x}^{(n)})$ and
@@
(3) $\hat{\bm{w}}^{\ell}=\underset{\bm{w}}{\operatorname{argmin}}\,\mathcal{L}(\bm{w})\implies\Delta\bm{w}^{\ell}=-\alpha\frac{\partial\mathcal{L}}{\partial\bm{w}^{\ell}},$
where the weights are updated iteratively using $\Delta\bm{w}^{\ell}$, with $\alpha$ the learning rate, and we use a hat to denote the optimized values. For better training efficiency, in practice one uses optimizers such as stochastic gradient descent (SGD) or Adam (Kingma and Ba, 2015).
@@ 2.1.2. Learning Algorithm
(5) $\Delta\bm{w}^{\ell}=\alpha\bm{\delta}^{\ell+1}\odot f^{\prime}(\bm{w}^{\ell}\bm{a}^{\ell})(\bm{a}^{\ell})^{T}\,,$
@@
Observe how errors are computed from the final to the first layer, explaining the name backpropagation of errors. Pseudocode is shown in alg. 1.
@@ 2.1.4. Testing Procedure
Finally, after a network has been trained, testing [^4] Note that ‘testing’ is often more generally called ‘inference’, to refer to use of a network after it has been trained. To prevent confusion with the inference phase during learning in PCNs, we refrain from using this term here. the network is done by using the activity rule (1) to obtain predicted outputs $\hat{\bm{y}}$ from datapoints $\bm{x}^{(n)}$.
@@ 2.2. Discriminative PCNs
Before explaining PCNs in detail we briefly mention a convention issue. The literature contains conflicting notions of direction in PCNs, which are worth discussing at the outset.
{mdframed}
[backgroundcolor=lightergray]
Direction in PCNs
There are two connected issues, the first being mathematical: should local predictions $\bm{\mu}^{\ell}$ be defined as $\bm{\mu}^{\ell}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$ (implied by fig. 1) or $\bm{\mu}^{\ell}=f(\bm{w}^{\ell-1}\bm{a}^{\ell-1})$ (as in this section)? The second issue is terminological: how should the words forward/backward be used?
@@
Typically in the neuroscientific PC literature (Rao and Ballard, 1999; Friston, 2008), predictions are $\bm{\mu}^{\ell}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$, and data is clamped to $\ell=0$ (fig. 1). This means that predictions go towards data (and errors go away from data). In this ‘PC convention’, the word forward is typically chosen to mean away from the data, which is the direction of errors (Rao and Ballard, 1999). However, in the supervised learning context, for a PCN to be equivalent to FNNs during testing, it turns out that one needs to have predictions go away from data. With the conventions above, this means one would have to clamp data to the highest layer $\ell=L$, which is confusing from the perspective of standard ML. This also leads to ambiguity as to whether ‘forward’ should mean towards increasing layer number, or away from the data.
@@
The alternative when comparing PCNs to FNNs is to keep data clamped to $\ell=0$ but change the direction of local prediction to $\bm{\mu}^{\ell}=f(\bm{w}^{\ell-1}\bm{a}^{\ell-1})$, as is done by many recent works (Millidge et al., 2022a; Rosenbaum, 2022; Salvatori et al., 2022b; Zahid et al., 2023a; Millidge et al., 2023a; Alonso et al., 2022; Millidge et al., 2023b; Innocenti et al., 2023; Song et al., 2024; Alonso et al., 2024; Byiringiro et al., 2022; Pinchetti et al., 2022). Forward then means away from the data (equivalently, towards higher layers) as in the PC convention above, except that this is now the direction of predictions. In other words, rather than swapping the data clamping from $\ell=0$ to $L$, we instead swap the directions of predictions and errors relative to fig. 1, so that – when comparing PCNs to FNNs specifically – predictions go forwards and errors go backwards.
@@
The disadvantage of this choice is that $\bm{\mu}^{\ell}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$ with data at $\ell=0$ is remains the more sensible convention for PCNs when seen as Bayesian/generative models (section 3) (Rao and Ballard, 1999; Salvatori et al., 2021a; Ororbia and Kifer, 2022; Ofner et al., 2022); it is only in the comparison with FNNs where some inversion (either of clamping or of directions) is required. Hence, in this work we use both conventions depending on the use of the network. We define discriminative PCNs as $\bm{\mu}^{\ell}=f(\bm{w}^{\ell-1}\bm{a}^{\ell-1})$ (for supervised learning), and generative PCNs as $\bm{\mu}^{\ell}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$ (for unsupervised learning). As for terminology, we can then unambiguously use ‘forward’ to mean away from the data in all cases, consistent with standard ML. One must simply bear in mind that for discriminative PCNs, ‘forward’ reflects the direction of predictions, while for generative PCNs it refers to the direction of errors.
@@ 2.2.1. Training Procedure
(not to be confused with the ‘global’ prediction $\hat{\bm{y}}$ of the network as a whole). [^5] Writing out bias, we here use $\bm{\mu}^{\ell}=f(\bm{w}^{\ell-1}\bm{a}^{\ell-1}+\bm{b}^{\ell-1})$, which is the most common convention in ML (Goodfellow et al., 2016; Prince, 2023). Most works in the PC literature however, define predictions as $\bm{\mu}^{\ell}=\bm{w}^{\ell-1}f(\bm{a}^{\ell-1})+\bm{b}^{\ell-1}$, whereas the main publicly available implementation uses yet another definition: $\bm{\mu}^{\ell}=f(\bm{w}^{\ell-1}\bm{a}^{\ell-1})+\bm{b}^{\ell-1}$ (Tschantz and Millidge, 2023). This choice of convention often does not make a difference for performance, but it is worth noting that it does change the form of the update rules. Also, the choice for the last layer is of great practical importance, since it determines the domain of possible outputs (e.g. a sigmoid or softmax is appropriate for classification but fatal for regression). The weights and biases are defined just like in FNNs. The discrepancy between actual and predicted activity is
@@
often called the energy function. [^6] In the probabilistic latent variable model perspective (section 3) this is the complete data log-likelihood. Section 3.4.3 further discusses the use of the term ‘energy’. Training now happens by setting the bottom layer $\bm{a}^{0}$ to the datapoint $\bm{x}^{(n)}$, like in FNNs, and additionally fixing the final layer $\bm{a}^{L}$ to the correct label $\bm{y}^{(n)}$. The activity rule is not defined as in (1); instead, the updated values of the hidden (unclamped) activity nodes (i.e. $\ell=1,...,L-1$) are the minimum of the energy function:
@@
This is called the inference phase, since we are inferring the appropriate activation values for hidden nodes, given the clamped nodes (cf. section 3 for a details of the probabilistic interpretation).
@@
$=-\gamma\left(\bm{\epsilon}^{\ell}-(\bm{w}^{\ell})^{T}\bm{\epsilon}^{\ell+1}\odot f^{\prime}(\bm{w}^{\ell}\bm{a}^{\ell})\right)$
@@
again using a hat for the values after minimizing, and a step-wise gradient descent updates which can be done using SGD or Adam. The fact that (11) is trained to convergence before the weight update is performed suggests that training a PCN is more computationally costly than FNNs. However, this added cost can be at least partially circumvented in practice, as we discuss below. Additionally, note that since the energy function is a sum over layers, the update rule does not require backpropagating errors. Instead, taking the derivative in (9) simply gives:
(13) $\Delta\bm{w}^{\ell}=\alpha\bm{\epsilon}^{l+1}\odot f^{\prime}(\bm{w}^{\ell}\hat{\bm{a}}^{\ell})(\hat{\bm{a}}^{\ell})^{T},$
i.e., the weight updates can be computed using the error and activation nodes in neighboring layers. This means PCNs’ computations are local both in space and time. We discuss the advantages of this property below.
@@ 2.2.2. Learning Algorithm
By virtue of the use of the energy function for both activation updates and weight updates, training PCNs is not done by backprop, but a different learning algorithm. Pseudocode is shown in alg. 2. This is sometimes called inference learning (IL) (Salvatori et al., 2021a; Alonso et al., 2022, 2024) or, rarely, prospective configuration (Song et al., 2024), but is often simply referred to as ‘predictive coding’ (Salvatori et al., 2022a, 2024b). Importantly, it is an instance of expectation maximization (EM) (see section 3). We choose the term IL in this work. Note that unlike BP, IL requires initializing hidden nodes before inference. This is often done using a ‘feedforward pass’, setting $\bm{a}^{\ell}(0)=\bm{\mu}^{\ell}$. This causes the error throughout the network to be zero except at the final layer. Other initializations can also be used (Frieder and Lukasiewicz, 2022).
@@
Importantly, recent work introduced incremental PC (Salvatori et al., 2024b) using the insight that EM admits the use of partial steps, meaning that the minimization of $E$ w.r.t. activations does not have to be complete for the algorithm to converge. Thus, a minor change in alg. 2 gives alg. 3, which we refer to as incremental IL in this work.
@@ 2.2.3. Structure
A PCN is typically visualized as in fig. 5. The grey connections are the predictions defined by (7) (similar to those in FNNs, see fig. 4), together with error computations, (8). The red connections are defined by (11).
@@ 2.2.4. Testing Procedure
A PCN can be tested by clamping the lowest layer $\bm{a}^{0}$ to the data, and computing updates for all the other layers, now including the final layer $\bm{a}^{L}$, using the activity rule (10). Interestingly, computing these updates is much simpler than during training. With the output nodes now unclamped, (10) has an analytical solution. One can show (Whittington and Bogacz, 2017; Song, 2021) (see appendix B.2) that with $\bm{a}^{0}=\bm{x}^{(n)}$, the minimum of $E$ is:
@@
In other words, during testing, a prediction of the network can be computed using a single pass through the network, without requiring gradient-based optimization of activation nodes of (11). In other words, we can even say that testing is equivalent in PCNs and FNNs, or that during testing the PCN ‘becomes’ an FNN in the sense of (2). As such, important results for neural networks such as universal approximation theorems (Prince, 2023) also hold for discriminative PCNs. For these networks, the difference w.r.t. FNNs is only in their learning algorithm. We emphasize that this holds only for discriminative PCNs, with the local prediction defined as in (7), and data clamped to $\bm{a}^{0}$. If this is changed, as will be done in later sections, then this equivalence no longer holds.
@@ 2.3.1. Training & Testing Procedures
Table 2 summarizes the important steps and computations for FNNs (with MSE loss) and discriminative PCNs. Notice the equality between the MSE loss and PCN’s energy function during testing, and their similarity during training. This is a result of fixing the final layer of the PCN $\bm{a}^{L}$ to the label $\bm{y}^{(n)}$, and labelling the prediction of the last layer $\bm{\mu}^{L}=f(\bm{w}^{L-1}\bm{a}^{L-1})$ as $\hat{\bm{y}}$. The term $\sum_{\ell=1}^{L-1}(\bm{\epsilon}^{\ell})^{2}$ is called the internal error (Millidge et al., 2023a) or residual error (Millidge et al., 2023b).
Table 2. Training & testing procedures compared: FNNs vs. discriminative PCNs. Key steps and corresponding equations are shown, and biases are omitted for clarity. ‘Fixed output’ refers to the final activity layer $\bm{a}^{L}$, clamped to data $\bm{y}^{(n)}$ during training (absent in FNNs). ‘Predicted output’ is the prediction in the final layer, $\bm{\mu}^{L}$, which becomes an output during testing, $\bm{\mu}^{L}=\hat{\bm{y}}$.